A collector feedback circuit uses an npn transistor with Vcc= 12V, Rc= 1.2V, RB = 47KΩ. Determine the collector voltage and the collector current if ß = 200.
1
Expert's answer
2020-10-07T06:33:02-0400

Since the circuit voltage is equal to the sum of the drop voltage and the collector voltage, "V_c=V_{cc}-R_c" , where "V_c" is the collector voltage, therefore

"V_c=12-1,2=10.8" V, since the current voltage falls to the base, therefore, according to the law of Oma "I_b=\frac {R_c}{R_b}" , and since "I_c=I_b\times β" ; "I_c=\frac {R_c}{R_b} \times β =\frac {1,2}{47\times10^3} \times 200 \approx 0.051" A.